# Average Winding Number for Determinantal Curves associated to 2-Matrix Models in the Class AIII

### v1

@@ 2.4 Main results
Expression (2.57) is not manifestly symmetric under the exchange of $a$ and $b$, nevertheless the underlying 2-matrix model (2.11) is symmetric and this invariance is preserved in a non-obvious way. More precisely, combining (2.37) and (2.58) allows to rewrite the deformed kernel $\widetilde{\mathrm{Q}}^{(N)}_{m}$ as a sum of eight terms for which the functions $a$ and $b$ clearly plays symmetrical roles. This expression is, however, too complicated to manipulate in further computations, so we deliberately retain the simpler asymmetric form which proves more natural and effective for the subsequent asymptotic analysis.
@@
$\mathbb{E}\left(\mathrm{Wind}_{N}\right)=\Bigg(\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{\nu\left(p\right)^{\dagger}\nu^{\prime}\left(p\right)}{\left\|\nu\left(p\right)\right\|^{2}}\mathrm{d}p\Bigg)N.$ (2.61)
@@
$\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{\nu\left(p\right)^{\dagger}\nu^{\prime}\left(p\right)}{\left\|\nu\left(p\right)\right\|^{2}}\mathrm{d}p=\frac{1}{2\pi}\oint_{\mathrm{S}^{1}}\mathcal{A}(p)\mathrm{d}p.$ (2.67)
@@
A thorough treatment of these notions: Aharonov-Anandan and Berry phases, holonomy, vector bundles and their realisations in quantum systems may be found in [26, Chap. 4-7,12].
@@
The average winding number associated with the determinantal curve of an additive two-matrix model where the two source matrices are independently drawn from a Muttalib-Borodin of Laguerre type ensemble admits, as $N$ goes to infinity, the following asymptotic expansion
$\begin{split}\mathbb{E}\left(\mathrm{Wind}_{N}\right)=&\Bigg(\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{\nu_{\gamma}\left(p\right)^{\dagger}\nu_{\gamma}^{\prime}\left(p\right)}{\left\|\nu_{\gamma}\left(p\right)\right\|^{2}}\mathrm{d}p\Bigg)N\\ &+\frac{\gamma-1-2\delta}{2}\Bigg(\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{\kappa^{\prime}\left(p\right)}{\kappa\left(p\right)}\frac{|a\left(p\right)|^{2\gamma}-|b\left(p\right)|^{2\gamma}}{|a\left(p\right)|^{2\gamma}+|b\left(p\right)|^{2\gamma}}\mathrm{d}p\Bigg)+o\left(1\right).\end{split}$ (2.70)
@@
$\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{\nu_{\gamma}\left(p\right)^{\dagger}\nu_{\gamma}^{\prime}\left(p\right)}{\left\|\nu_{\gamma}\left(p\right)\right\|^{2}}\mathrm{d}p=\frac{1}{2\pi i}\oint_{\rm{S}^{1}}\frac{|a\left(p\right)|^{2\gamma-2}a^{\prime}\left(p\right)\overline{a\left(p\right)}+|b\left(p\right)|^{2\gamma-2}b^{\prime}\left(p\right)\overline{b\left(p\right)}}{|a\left(p\right)|^{2\gamma}+|b\left(p\right)|^{2\gamma}}\mathrm{d}p.$ (2.71)
@@
The integral appearing in the sub-leading order correction term must be interpreted in the sense of a Cauchy principal value integral whenever $a$ or $b$ vanishes. Since we assume that both functions have only isolated zeroes on $\mathrm{S}^{1}$, their logarithmic derivatives possess at most simple poles. Consequently, the Cauchy principal value ensures that these singularities remain integrable.
It should be noted that $\left\|\nu_{\gamma}\left(p\right)\right\|$ is non-zero for all $p\in\mathrm{S}^{1}$, since by assumption the parameter functions $a$ and $b$ never vanish simultaneously. The term (2.71) remains an Aharonov-Anandan phase angle, now associated with the deformed vector field $\nu_{\gamma}$. The additional constant term is genuinely new and does not occur in the Gaussian case $\gamma=1$, cf. (2.61), indeed it vanishes in the joint limit $\gamma\to 1$ and $\delta\to 0$. Interestingly, the tail behaviour and the repulsion from the origin can compensate each other in the sub-leading order contribution when the parameters are tuned such that $\gamma-1=2\delta$.
More interestingly, when neither $a$ nor $b$ vanishes on $\mathrm{S}^{1}$, the sub-leading order term admits a geometric interpretation as a deformed winding number of the curve traced by $\kappa=a/b$ around the origin. The deformation is controlled by a weight that measures the local asymmetry between $a$ and $b$. The contribution takes values in [1,1][-1,1], vanishes wherever $|a(p)|=|b(p)|$ and approaches unity when $\left|\ln(\kappa(p))\right|$ is large, that is, when one component of the vector $\nu$ dominates in magnitude.
@@
In the particular case where $r_{1}=r_{2}$ the average winding number becomes a half-integer which is noteworthy, although this behaviour does not necessarily holds in the general case.
@@
In the previous example it appears that the leading order may vanish while the sub-leading order term persists, for example setting $r_{1}=2^{1/2\gamma},\,r_{2}=1$ and 2n=m2n=-m yields an average winding number asymptotically equal to $n(\gamma-1-2\delta)/2$ as $N$ goes to infinity.
@@ 2.5 The Ginibre random field
We now analyse the partition function (2.6) in the case $m=1$ and its associated 1-point correlation function (2.8) beyond the additive two-matrix model framework and consider a matrix-valued Gaussian random field over the unit circle $\mathrm{S}^{1}$, whose entries are independent and identically distributed complex Gaussian random fields.
A complex Gaussian random field is uniquely determined by its mean, covariance and pseudo-covariance kernels. In the present case the mean function is equally zero and the covariance/pseudo-covariance kernels are given by
@@
for all indices $i,j,k,\ell\in\left[1,N\right]$ and points $p,q\in\rm{S}^{1}$. The covariance kernel $\mathrm{C}$ is Hermitian and positive-definite, in particular:
@@
At first glance the choice of a null pseudo-covariance function might seem restrictive but it actually reflects the independence of the real and imaginary part of each entry in the definition of the Ginibre Unitary Ensemble. This setting naturally generalizes the construction studied in [11, 10], which may be recovered by specifying
$\mathrm{C}(p,q)=\nu(q)^{\dagger}\mathrm{J}\nu(p).$ (2.76)
A random field $K$ as in (2.74) may be referred to as a complex Ginibre random field (of size $N$ with covariance kernel $\mathrm{C}$ over $\mathrm{S}^{1}$). To analyse functionals of this random field, we sample it at finitely many points on the base manifold $\mathrm{S}^{1}$, say $n\geq 1$. The result joint distribution of the random matrices $K(p_{1}),\ldots,K(p_{n})$ admits a finite-dimensional representation in terms of $n$ independent GinUE-drawn random matrices. This construction is analogous to a Karhunen-Loève expansion, see for example [27, Sec. 3.2].
@@
Making use of the entry-wise independence of the random field $K$, we start by fixing two indices $i,j$ in $[1,N]$ and defining the random vector
@@
Its covariance matrix is $\mathrm{C}_{n}(\mathbf{p})=(\mathrm{C}(p_{k},p_{\ell}))_{1\leq k,\ell\leq n}$ which is Hermitian and positive. Therefore the Cholesky decomposition ensures the existence of a unique lower-triangular matrix $\mathrm{L}(\mathbf{p})$ with positive diagonal entries such that $\mathrm{C}_{n}(\mathbf{p})=\mathrm{L}(\mathbf{p})\mathrm{L}(\mathbf{p})^{\dagger}$.
@@
$\begin{split}\mathrm{L}_{1,1}(\mathbf{p})=&\sqrt{{\rm C}(p_{1},p_{1})},\\ \mathrm{L}_{i,i}(\mathbf{p})=&\sqrt{{\rm C}(p_{i},p_{i})-\sum_{k=1}^{i-1}|\mathrm{L}_{i,k}(\mathbf{p})|^{2}},\\ \mathrm{L}_{i,j}(\mathbf{p})=&\mathrm{L}_{i,i}(\mathbf{p})^{-1}\left({\rm C}(p_{i},p_{j})-\sum_{k=1}^{j-1}\overline{\mathrm{L}_{i,k}(\mathbf{p})}\mathrm{L}_{j,k}(\mathbf{p})\right),\end{split}$ (2.79)
coupled to the fact that $\mathrm{C}$ being positive-definite, all the principal minors of $\mathrm{C}_{n}(\mathbf{p})$ are positive. Therefore the square-roots as well as the inverses are well-defined. ∎
@@
$\left(\sqrt{{\rm C}(q,q)}G_{1},\frac{{\rm C}(p,q)G_{1}+\sqrt{{\rm C}(q,q){\rm C}(p,p)-|{\rm C}(p,q)|^{2}}\,G_{2}}{\sqrt{{\rm C}(q,q)}}\right)$
@@
$\mathscr{Z}_{1}^{\left(N\right)}\left(p,q\right)={\rm C}(q,q)^{-N}\mathbb{E}\Big(\det({\rm C}(p,q)I_{N}+\sqrt{{\rm C}(q,q){\rm C}(p,p)-|{\rm C}(p,q)|^{2}}\,G_{1}^{-1}G_{2})\Big).$
The distribution of the matrix $Y=G_{1}^{-1}G_{2}$ coincides with that of the complex spherical ensemble. In particular its law is $\mathrm{U}(1)$-invariant in the following sense: for all $\theta\in[0,2\pi[$, $Y\overset{d}{=}e^{i\theta}Y$. It implies that only the scalar term $\mathrm{C}(p,q)$ contributes in the expectation, leading to the first identity (2.80). The second identity follows directly by plugging (2.80) into (2.3) and (2.8). ∎
@@
Following [11], the partition function $\mathscr{Z}^{\left(N\right)}_{m_{1},m_{2}}$, see (2.51), can be recast into the form
@@
It is then possible to exploit the closure of Pólya ensembles under matrix multiplication and inversion, thereby reducing the 2-matrix model to an effective 1-matrix model. We recall that if $K_{1}$ and $K_{2}$ are two random matrices independently drawn from a Pólya ensemble with weight $\omega$ then the generalized ratio $K_{1}^{-1}K_{2}$ is drawn from a Pólya ensemble with weight $\widehat{\omega}$ defined at (2.34). Furthermore, the random matrix $K_{1}^{-1}K_{2}$ admits the following complex eigenvalues jpdf
@@
with $\mathrm{d}^{2}\mathbf{z}=\prod\limits_{j=1}^{N}\mathrm{d}\Re\left(z_{j}\right)\mathrm{d}\Im\left(z_{j}\right)$. Such integrals are known and have been computed in a general setting in [28, Eq. (4.18)].
@@
$\begin{split}&\widetilde{\mathscr{Z}}_{\frac{n_{1}/m_{1}}{n_{2}/m_{2}}}\left(\textbf{p},\textbf{q},\widetilde{\textbf{p}},\widetilde{\textbf{q}}\right)\\ =&\int_{\mathbb{C}^{N}}\left|\Delta\left(\mathbf{z}\right)\right|^{2}\prod\limits_{j=1}^{N}\widehat{\omega}\left(\left|z_{j}\right|^{2}\right)\frac{\prod\limits_{k_{1}=1}^{m_{1}}\Big(-\kappa\left(p_{k_{1}}\right)-z_{j}\Big)}{\prod\limits_{\ell_{1}=1}^{n_{1}}\Big(-\kappa\left(q_{\ell_{1}}\right)-z_{j}\Big)}\frac{\prod\limits_{k_{2}=1}^{m_{2}}\Big(-\overline{\kappa\left(\widetilde{p}_{k_{2}}\right)}-\overline{z_{j}}\Big)}{\prod\limits_{\ell_{2}=1}^{n_{2}}\Big(-\overline{\kappa\left(\widetilde{q}_{\ell_{2}}\right)}-\overline{z_{j}}\Big)}\mathrm{d}^{2}\mathbf{z},\end{split}$ (3.4)
where $m_{1},m_{2},n_{1},n_{2}$ are four integers such that $n_{1}-m_{1}+N=n_{2}-m_{2}+N\geq 0$. Our goal is to compute this integral in the sub-case $m_{1}=n_{1}$ and $m_{2}=n_{2}$, we, however, define it in greater generality as kernels related to the partition functions $\widetilde{\mathscr{Z}}_{\frac{0/1}{0/1}}$ and $\widetilde{\mathscr{Z}}_{\frac{1/0}{1/0}}$ will appear later on. To that end, we also introduce the following auxiliary matrices which are expressed using these functions,
@@
Then, the application of the identity [28, Eq. (4.35)] to (3.4) yields the compact form
@@
$=$ $\frac{1}{\mathcal{M}[\widehat{\omega}](N)C^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]}\Bigg[\mathcal{S}[\widehat{\omega}\sigma_{r,R}]\left(\kappa(q_{i})\overline{\kappa(\tilde{q}_{j})}\right)$ (3.14)
$-\sum_{k=1}^{N-1}\frac{\mathcal{M}[\widehat{\omega}](k,\left|\kappa(q_{i})\right|^{2})\mathcal{M}[\widehat{\omega}](k,\left|\kappa(\widetilde{q}_{j})\right|^{2})}{\mathcal{M}[\widehat{\omega}](k)(\kappa(q_{i})\overline{\kappa(\tilde{q}_{j})})^{k}}\Bigg],$ (3.15)
@@
The kernel (3.12) is, actually, the pre-kernel of the $k$-point correlation function for the eigenvalues of the Pólya ensemble associated to the weight function $\widehat{\omega}$, see [12, Lemma 4.1]. In particular,
$\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{0/1}{0/1}}\left(p_{i},\widetilde{p}_{j}\right)=\frac{\pi\mathcal{M}[\widehat{\omega}](N+1)}{C^{\left(N\right)}_{\mathrm{EV}}}\frac{K^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]\left(\kappa(p_{i}),\kappa(\widetilde{p}_{j})\right)}{\sqrt{\widehat{\omega}\left(\left|\kappa(p_{i})\right|^{2}\right)\widehat{\omega}\left(\left|\kappa(\widetilde{p}_{j})\right|^{2}\right)}},$ (3.16)
which reduced to a relation to the level density when $\widetilde{p}_{j},p_{i}\to p$, i.e.,
$\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{0/1}{0/1}}\left(p,p\right)=\frac{\pi\mathcal{M}[\widehat{\omega}](N+1)}{C^{\left(N\right)}_{\mathrm{EV}}}\frac{\rho^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]\left(\left|\kappa\left(p\right)\right|^{2}\right)}{\widehat{\omega}\left(\left|\kappa\left(p\right)\right|^{2}\right)}.$ (3.17)
@@
$\mathrm{T}^{\left(N\right)}_{m_{1},m_{2}}\left[\widehat{\omega}\right]\left(\mathbf{q},\widetilde{\mathbf{q}}\right)=\left(C_{\rm EV}^{(N)}[\widehat{\omega}]\frac{\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{1/0}{1/0}}\left(q_{i},\widetilde{q}_{j}\right)}{(b(q_{i})\overline{b(q_{j})})^{N}}\right)_{\begin{subarray}{c}1\leq i\leq m_{1}\\ 1\leq j\leq m_{2}\end{subarray}}.$ (3.19)
@@
Theorem 2.8 then follows directly by setting $m_{1}=m$ and $m_{2}=0$. Under the assumptions of definition (2.51), the three kernel functions are well-defined at all points $(q_{i},\widetilde{q}_{j})$ such that $b(q_{i})=0$ or $b(\widetilde{q}_{j})=0$ which guarantees the existence of the integral representation.

### v2

@@ 2.4 Main results
Expression (2.56) is not manifestly symmetric under the exchange of $a$ and $b$, nevertheless the underlying 2-matrix model (2.11) is symmetric and this invariance is preserved in a non-obvious way. More precisely, combining (2.37) and (2.57) allows to rewrite the deformed kernel $\widetilde{\mathrm{Q}}^{(N)}_{m}$ as a sum of eight terms for which the functions $a$ and $b$ clearly play symmetrical roles. This expression is, however, too complicated to manipulate in further computations, so we deliberately retain the simpler asymmetric form which proves more natural and effective for the subsequent asymptotic analysis.
@@
$\mathbb{E}\left(\mathrm{Wind}_{N}\right)=\Bigg(\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{\nu\left(p\right)^{\dagger}\nu^{\prime}\left(p\right)}{\left\|\nu\left(p\right)\right\|^{2}}\mathrm{d}p\Bigg)N.$ (2.60)
@@
$\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{\nu\left(p\right)^{\dagger}\nu^{\prime}\left(p\right)}{\left\|\nu\left(p\right)\right\|^{2}}\mathrm{d}p=\frac{1}{2\pi}\oint_{\mathrm{S}^{1}}\mathcal{A}(p)\mathrm{d}p.$ (2.66)
@@
A thorough treatment of these notions: Aharonov-Anandan and Berry phases, holonomy, vector bundles and their realisations in quantum systems may be found in [BohmGeometricPhaseQuantum2003, Chap. 4-7,12].
@@
The average winding number associated with the determinantal curve of an additive $2$-matrix model where the two source matrices are independently drawn from a Muttalib-Borodin of Laguerre type ensemble admits, as $N$ goes to infinity, the following asymptotic expansion
$\begin{split}\mathbb{E}\left(\mathrm{Wind}_{N}\right)=&\Bigg(\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{\nu_{\gamma}\left(p\right)^{\dagger}\nu_{\gamma}^{\prime}\left(p\right)}{\left\|\nu_{\gamma}\left(p\right)\right\|^{2}}\mathrm{d}p\Bigg)N\\ &+\frac{\gamma-1-2\delta}{2}\Bigg(\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{\kappa^{\prime}\left(p\right)}{\kappa\left(p\right)}\frac{|a\left(p\right)|^{2\gamma}-|b\left(p\right)|^{2\gamma}}{|a\left(p\right)|^{2\gamma}+|b\left(p\right)|^{2\gamma}}\mathrm{d}p\Bigg)+o\left(1\right).\end{split}$ (2.69)
@@
$\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{\nu_{\gamma}\left(p\right)^{\dagger}\nu_{\gamma}^{\prime}\left(p\right)}{\left\|\nu_{\gamma}\left(p\right)\right\|^{2}}\mathrm{d}p=\frac{1}{2\pi i}\oint_{\mathrm{S}^{1}}\frac{|a\left(p\right)|^{2\gamma-2}a^{\prime}\left(p\right)\overline{a\left(p\right)}+|b\left(p\right)|^{2\gamma-2}b^{\prime}\left(p\right)\overline{b\left(p\right)}}{|a\left(p\right)|^{2\gamma}+|b\left(p\right)|^{2\gamma}}\mathrm{d}p.$ (2.70)
@@
The integral appearing in the subleading order correction term must be interpreted in the sense of a Cauchy principal value integral whenever $a$ or $b$ vanishes. Since we assume that both functions have only isolated zeroes on $\mathrm{S}^{1}$, their logarithmic derivatives possess at most simple poles. Consequently, the Cauchy principal value ensures that these singularities remain integrable.
It should be noted that $\left\|\nu_{\gamma}\left(p\right)\right\|$ is non-zero for all $p\in\mathrm{S}^{1}$, since by assumption the parameter functions $a$ and $b$ never vanish simultaneously. The term (2.70) remains an Aharonov-Anandan phase angle, now associated with the deformed vector field $\nu_{\gamma}$. The additional constant term is genuinely new and does not occur in the Gaussian case $\gamma=1$, cf. (2.60), indeed it vanishes in the joint limit $\gamma\to 1$ and $\delta\to 0$. Interestingly, the tail behaviour and the repulsion from the origin can compensate each other in the subleading order contribution when the parameters are tuned such that $\gamma-1=2\delta$.
More interestingly, when neither $a$ nor $b$ vanishes on $\mathrm{S}^{1}$, the subleading order term admits a geometric interpretation as a deformed winding number of the curve traced by $\kappa=a/b$ around the origin. The deformation is controlled by a weight that measures the local asymmetry between $a$ and $b$. The contribution takes values in [1,1][-1,1], vanishes wherever $|a(p)|=|b(p)|$ and approaches unity when $\left|\ln(\kappa(p))\right|$ is large, that is, when one component of the vector $\nu$ dominates in magnitude.
@@
In the particular case where $r_{1}=r_{2}$, the average winding number becomes a half-integer which is noteworthy, although this behaviour does not necessarily hold in the general case.
@@
In the previous example it appears that the leading order may vanish while the subleading order term persists, for example setting $r_{1}=2^{1/2\gamma},\,r_{2}=1$ and 2n=m2n=-m yields an average winding number asymptotically equal to $n(\gamma-1-2\delta)/2$ as $N$ goes to infinity.
@@ 2.5 The Ginibre random field
We now analyse the partition function (2.6) in the case $m=1$ and its associated 1-point correlation function (2.8) beyond the additive $2$-matrix model framework and consider a matrix-valued Gaussian random field over the unit circle $\mathrm{S}^{1}$, whose entries are independent and identically distributed complex Gaussian random fields.
A complex Gaussian random field is uniquely determined by its mean, covariance and pseudo-covariance kernels. In the present case the mean function is identically zero and the covariance/pseudo-covariance kernels are given by
@@
for all indices $i,j,k,\ell\in\left[1,N\right]$ and points $p,q\in\mathrm{S}^{1}$. The covariance kernel $\mathrm{C}$ is Hermitian and positive-definite, in particular:
@@
At first glance the choice of a null pseudo-covariance function might seem restrictive but it actually reflects the independence of the real and imaginary parts of each entry in the definition of the Ginibre Unitary Ensemble. This setting naturally generalizes the construction studied in [hahnWindingNumberStatistics2023, hahnUniversalCorrelations2024], which may be recovered by specifying
$\mathrm{C}(p,q)=\nu(q)^{\dagger}\nu(p).$ (2.75)
A random field $K$ as in (2.73) may be referred to as a complex Ginibre random field (of size $N$ with covariance kernel $\mathrm{C}$ over $\mathrm{S}^{1}$). To analyse functionals of this random field, we sample it at finitely many points on the base manifold $\mathrm{S}^{1}$, say $n\geq 1$. The result joint distribution of the random matrices $K(p_{1}),\ldots,K(p_{n})$ admits a finite-dimensional representation in terms of $n$ independent GinUE-drawn random matrices. This construction is analogous to a Karhunen-Loève expansion, see for example [taylorRandomFieldsGeometry2007, Sec. 3.2].
@@
Making use of the entrywise independence of the random field $K$, we start by fixing two indices $i,j$ in $[1,n]$ and defining the random vector
@@
Its covariance matrix is $\mathrm{C}_{n}(\mathbf{p})=(\mathrm{C}(p_{k},p_{\ell}))_{1\leq k,\ell\leq n}$ which is Hermitian and positive definite. Therefore the Cholesky decomposition ensures the existence of a unique lower-triangular matrix $\mathrm{L}(\mathbf{p})$ with positive diagonal entries such that $\mathrm{C}_{n}(\mathbf{p})=\mathrm{L}(\mathbf{p})\mathrm{L}(\mathbf{p})^{\dagger}$.
@@
$\begin{split}\mathrm{L}_{1,1}(\mathbf{p})=&\sqrt{\mathrm{C}(p_{1},p_{1})},\\ \mathrm{L}_{i,i}(\mathbf{p})=&\sqrt{\mathrm{C}(p_{i},p_{i})-\sum_{k=1}^{i-1}|\mathrm{L}_{i,k}(\mathbf{p})|^{2}},\\ \mathrm{L}_{i,j}(\mathbf{p})=&\mathrm{L}_{i,i}(\mathbf{p})^{-1}\left(\mathrm{C}(p_{i},p_{j})-\sum_{k=1}^{j-1}\overline{\mathrm{L}_{i,k}(\mathbf{p})}\mathrm{L}_{j,k}(\mathbf{p})\right),\end{split}$ (2.78)
coupled to the fact that $\mathrm{C}$ being positive-definite: all the principal minors of $\mathrm{C}_{n}(\mathbf{p})$ are positive. Therefore the square-roots as well as the inverses are well-defined. ∎
@@
$\left(\sqrt{\mathrm{C}(q,q)}G_{1},\frac{\mathrm{C}(p,q)G_{1}+\sqrt{\mathrm{C}(q,q)\mathrm{C}(p,p)-|\mathrm{C}(p,q)|^{2}}\,G_{2}}{\sqrt{\mathrm{C}(q,q)}}\right)$
@@
$\mathscr{Z}_{1}^{\left(N\right)}\left(p,q\right)=\mathrm{C}(q,q)^{-N}\mathbb{E}\Big(\det(\mathrm{C}(p,q)I_{N}+\sqrt{\mathrm{C}(q,q)\mathrm{C}(p,p)-|\mathrm{C}(p,q)|^{2}}\,G_{1}^{-1}G_{2})\Big).$
The distribution of the matrix $Y=G_{1}^{-1}G_{2}$ coincides with that of the complex spherical ensemble. In particular, its law is $\mathrm{U}(1)$-invariant in the following sense: for all $\theta\in[0,2\pi[$, $Y\overset{d}{=}e^{i\theta}Y$. This implies that only the scalar term $\mathrm{C}(p,q)$ contributes in the expectation, leading to the first identity (2.79). The second identity follows directly by plugging (2.79) into (2.3) and (2.8). ∎
@@
Following [hahnWindingNumberStatistics2023], the partition function $\mathscr{Z}^{\left(N\right)}_{m_{1},m_{2}}$, see (2.50), can be recast into the form
@@
It is then possible to exploit the closure of Pólya ensembles under matrix multiplication and inversion, thereby reducing the 2-matrix model to an effective 1-matrix model. We recall that if $K_{1}$ and $K_{2}$ are two random matrices independently drawn from a Pólya ensemble with weight $\omega$ then the generalized ratio $K_{1}^{-1}K_{2}$ is drawn from a Pólya ensemble with weight $\widehat{\omega}$ defined at (2.34). Furthermore, the random matrix $K_{1}^{-1}K_{2}$ admits the following jpdf for complex eigenvalues
@@
with $\mathrm{d}^{2}\mathbf{z}=\prod\limits_{j=1}^{N}\mathrm{d}\Re\left(z_{j}\right)\mathrm{d}\Im\left(z_{j}\right)$. Such integrals are known and have been computed in a general setting in [kieburgDerivationDeterminantalStructures2010, Eq. (4.18)].
@@
$\begin{split}&\widetilde{\mathscr{Z}}_{\frac{n_{1}/m_{1}}{n_{2}/m_{2}}}^{\left(N\right)}\left(\textbf{p},\textbf{q},\widetilde{\textbf{p}},\widetilde{\textbf{q}}\right)\\ =&\int_{\mathbb{C}^{N}}\left|\Delta_{N}\left(\mathbf{z}\right)\right|^{2}\prod\limits_{j=1}^{N}\widehat{\omega}\left(\left|z_{j}\right|^{2}\right)\frac{\prod\limits_{k_{1}=1}^{m_{1}}\Big(-\kappa\left(p_{k_{1}}\right)-z_{j}\Big)}{\prod\limits_{\ell_{1}=1}^{n_{1}}\Big(-\kappa\left(q_{\ell_{1}}\right)-z_{j}\Big)}\frac{\prod\limits_{k_{2}=1}^{m_{2}}\Big(-\overline{\kappa\left(\widetilde{p}_{k_{2}}\right)}-\overline{z_{j}}\Big)}{\prod\limits_{\ell_{2}=1}^{n_{2}}\Big(-\overline{\kappa\left(\widetilde{q}_{\ell_{2}}\right)}-\overline{z_{j}}\Big)}\mathrm{d}^{2}\mathbf{z},\end{split}$ (3.4)
where $m_{1},m_{2},n_{1},n_{2}$ are four integers such that $n_{1}-m_{1}+N=n_{2}-m_{2}+N\geq 0$. Our goal is to compute this integral in the sub-case $m_{1}=n_{1}$ and $m_{2}=n_{2}$. We, however, define it in greater generality as kernels related to the partition functions $\widetilde{\mathscr{Z}}_{\frac{0/1}{0/1}}$ and $\widetilde{\mathscr{Z}}_{\frac{1/0}{1/0}}$ will appear later on. To that end, we also introduce the following auxiliary matrices which are expressed using these functions,
@@
Then, the application of the identity [kieburgDerivationDeterminantalStructures2010, Eq. (4.35)] to (3.4) yields the compact form
@@
$=$ $\frac{1}{\mathcal{M}[\widehat{\omega}](N)C^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]}\Bigg[\mathcal{S}[\widehat{\omega}\sigma_{r,R}]\left(\kappa(q_{i})\overline{\kappa(\widetilde{q}_{j})}\right)$ (3.14)
$-\sum_{k=1}^{N-1}\frac{\mathcal{M}[\widehat{\omega}](k,\left|\kappa(q_{i})\right|^{2})\mathcal{M}[\widehat{\omega}](k,\left|\kappa(\widetilde{q}_{j})\right|^{2})}{\mathcal{M}[\widehat{\omega}](k)(\kappa(q_{i})\overline{\kappa(\widetilde{q}_{j})})^{k}}\Bigg],$ (3.15)
@@
The kernel (3.12) is, actually, the pre-kernel of the $k$-point correlation function for the eigenvalues of the Pólya ensemble associated to the weight function $\widehat{\omega}$, see [kieburgExactRelationSingular2016, Lemma 4.1]. In particular,
$\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{0/1}{0/1}}\left(p_{i},\widetilde{p}_{j}\right)=\frac{\pi\mathcal{M}[\widehat{\omega}](N+1)}{C^{\left(N\right)}_{\mathrm{EV}}[\widehat{\omega}]}\frac{K^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]\left(\kappa(p_{i}),\kappa(\widetilde{p}_{j})\right)}{\sqrt{\widehat{\omega}\left(\left|\kappa(p_{i})\right|^{2}\right)\widehat{\omega}\left(\left|\kappa(\widetilde{p}_{j})\right|^{2}\right)}},$ (3.16)
which reduces to a relation to the level density when $\widetilde{p}_{j},p_{i}\to p$, i.e.,
$\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{0/1}{0/1}}\left(p,p\right)=\frac{\pi\mathcal{M}[\widehat{\omega}](N+1)}{C^{\left(N\right)}_{\mathrm{EV}}[\widehat{\omega}]}\frac{\rho^{\left(N\right)}_{\mathrm{EV}}\left[\widehat{\omega}\right]\left(\left|\kappa\left(p\right)\right|^{2}\right)}{\widehat{\omega}\left(\left|\kappa\left(p\right)\right|^{2}\right)}.$ (3.17)
@@
$\mathrm{T}^{\left(N\right)}_{m_{1},m_{2}}\left[\widehat{\omega}\right]\left(\mathbf{q},\widetilde{\mathbf{q}}\right)=\left(C_{\rm EV}^{(N)}[\widehat{\omega}]\frac{\widetilde{\mathscr{Z}}^{\left(N\right)}_{\frac{1/0}{1/0}}\left(q_{i},\widetilde{q}_{j}\right)}{(b(q_{i})\overline{b(\widetilde{q}_{j})})^{N}}\right)_{\begin{subarray}{c}1\leq i\leq m_{1}\\ 1\leq j\leq m_{2}\end{subarray}}.$ (3.19)
@@
Theorem 2.8 then follows directly by setting $m_{1}=m$ and $m_{2}=0$. Under the assumptions of Definition (2.50), the three kernel functions are well-defined at all points $(q_{i},\widetilde{q}_{j})$ such that $b(q_{i})=0$ or $b(\widetilde{q}_{j})=0$, which guarantees the existence of the integral representation.